OpenStudy (anonymous):

Simplify the expression -3 over 8 - 1 over 10 A. -3 over 8 B. 19 over 40 C. 2 over 9 D. -19 over 40

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Thank you. Gave you a medal.

OpenStudy (unklerhaukus):

\[{-3 \over 8} - {1 \over 10}\\=\frac{-3}8\times\frac{10}{10}-\frac88\frac1{10}\\=\frac{-3\times10-8\times1}{8\times10}\\=\]